# DragPoser: Motion Reconstruction from Variable Sparse Tracking Signals via Latent Space Optimization

### v1

@@ 5. Experiments and Evaluation
For evaluation, we trained DragPoser using the DanceDB dataset (Aristidou et al., 2019) and evaluated with the HUMAND4D (Chatzitofis et al., 2020) and SOMA (Ghorbani and Black, 2021) datasets, both part of the AMASS collection (Mahmood et al., 2019). Our choice of AMASS was guided by its public availability and its use in previous works. Additionally, for generating other images and supplemental material, we employed the SparsePoser database (Ponton et al., 2023). This database is particularly advantageous for VR applications, offering diverse common motions and reducing incorrect poses through the use of IMU-based motion capture compared to the optical systems used in AMASS datasets.
@@
We implemented DragPoser using PyTorch (Paszke et al., 2019) and optimized the networks with the AdamW optimizer (Loshchilov and Hutter, 2019). The system was trained and evaluated on a PC configured with an Intel Core i7-12700k CPU, 32GB of RAM, and an NVIDIA GeForce RTX 3090 GPU. The Pose Autoencoder was trained with a batch size of 64 and a learning rate of $10^{-4}$ over approximately one day; we set the latent space dimension to $L=24$ and the number of joints to $J=22$, resulting in about 168 thousand parameters. For the Temporal Predictor, we trained it with a batch size of 512 and a learning rate of $10^{-3}$ for approximately thirty minutes. The Transformer hyperparameters were set to four heads, three layers each for the encoder and decoder, a feedforward dimension of 2048, and a feature dimension of 48, totaling approximately 1.2 million parameters. During training, we used $W_{p}=16$ and $W_{f}=16$ and ran the Pose Autoencoder’s Encoder to generate the training database for the Temporal Predictor. Additionally, we empirically found that introducing random noise to the four limbs aids in regularizing the Temporal Predictor. Thus, in every training iteration, we introduce Gaussian noise to the latent vector, with mean and standard deviation values for each dimension derived from the dataset, applied to each limb with a 10% probability. We initially experimented with adding noise to arbitrary joints in the full skeleton, but we found that targeting individual limbs was more effective. First we train the Pose Autoencoder with a weighted combination of all the losses presented $\mathcal{L}=\lambda_{Q}\mathcal{L}_{Q}+\lambda_{FK}\mathcal{L}_{FK}+\lambda_{KLD}\mathcal{L}_{KLD}+\lambda_{C}\mathcal{L}_{C}$ where $\lambda_{Q}=1$, $\lambda_{FK}=100$, $\lambda_{KLD}=0.001$ and $\lambda_{C}=1$. Subsequently, we train the Temporal Predictor with a standard MSE loss.
Performance
DragPoser can be used for both offline and real-time motion reconstruction applications. We measured the average execution times of the different components. Note that our code is executed in Python with no special optimizations. On average, the Temporal Predictor forward pass takes around 7 ms per frame. The other components, which run each optimization iteration, have the following performance: the Decoder’s forward pass averages 0.63 ms, the constraints (losses) computation takes around 4.59 ms (mainly because of the FK), and the backpropagation through the Decoder approximately 6.91 ms. These three processes are computed on the CPU due to the relatively lightweight Decoder model. For real-time applications, we typically limit the number of iterations to 1 per frame, achieving a processing time of about 19 ms per frame. In contrast, we allow up to 10 iterations per frame for offline pose processing, which extends the processing time to roughly a maximum of 128 ms per frame. This flexibility enables DragPoser to adapt to the performance requirements of various application scenarios.
5.1. Comparison
We compare DragPoser with two state-of-the-art deep-learning-based methods, AvatarPoser (Jiang et al., 2022a) and SparsePoser (Ponton et al., 2023), as well as FinalIK (RootMotion, 2017), a state-of-the-art IK method for animating full-body characters. These methods were selected due to their public availability and their representation of different sensor configurations and approaches. AvatarPoser utilizes a Transformer encoder to generate full-body poses from sparse input (three or four sensors) and refines the arms with an optimization-based IK. Notably, it excels in producing high-quality poses in scenarios with high ambiguity and limited sensors. We extended the method, referred to here as AvatarPoser+, by adding two additional sensors on the feet, allowing a fair comparison in six sensor setups. In addition, we compare with SparsePoser, which is the state-of-the-art deep-learning-based method in terms of pose quality and end-effector accuracy for six 6-DoF sensors. Their approach is fixed to six input sensors and uses an autoencoder and learned IK networks for full-body pose reconstruction. Both AvatarPoser and SparsePoser were re-trained using the DanceDB dataset following the training procedure of their official implementations. For AvatarPoser, we incorporated DanceDB into its pre-existing training database. Finally, considering the findings by Ponton et al. (2023) that FinalIK exhibits superior end-effector accuracy compared to deep learning-based methods, it is employed as a baseline for evaluating hard constraints.
Method Sensors Joint Error End-Effector Error
@@ 5.1. Comparison
AvatarPoser+ 6 7.66(7.60) 11.3(10.5) 13.4(28.5) 11.2(7.93)
FinalIK 6 3.31(4.64) 11.5(19.2) 6.18(21.2) 1.16(1.22)
SparsePoser 6 2.81(2.66) 5.78(5.85) 10.8(15.0) 3.43(2.53)
Ours 6 2.18(2.70) 7.66(8.71) 10.7(15.1) 1.02(0.73)
AvatarPoser 4 9.24(10.1) 12.4(11.7) 16.3(31.7) 9.62(6.51)
Ours 4 7.86(13.1) 10.6(12.4) 20.0(41.5) 1.12(0.72)
AvatarPoser 3 11.0(10.8) 14.3(12.8) 20.1(39.5) 12.3(7.51)
Ours Head+Hands 3 15.5(15.0) 22.0(22.3) 34.0(50.2) 16.1(11.4)
Ours Hip+Hands 3 9.95(13.7) 14.1(15.6) 23.1(36.2) 1.77(1.31)
Table 1. Accuracy of the reconstructed poses compared against state-of-the-art methods, for diverse sensor configurations. We report the mean and standard deviation (within parenthesis) of joint errors (position, rotation, and velocity) as well as the end-effector positional error, all of them with respect to ground truth data. Please refer to the text for details about training and test datasets. When using 6 sensors (placed at the hip, head, hands, and feet), we could compare against AvatarPoser+ (i.e., (Jiang et al., 2022b) extended to use also the two sensors on the feet), FinalIK (RootMotion, 2017), and SparsePoser (Ponton et al., 2023). For the 4-sensor configuration (hip, head, and hands), we compare against AvatarPoser; FinalIK is not considered since it requires all end-effectors, and SparsePoser’s architecture is designed for a fixed 6-sensor configuration. Concerning the 3-sensor setup, we tested our approach with a head+hands configuration and with a hip+hands one, and compared it against AvatarPoser.
@@
In the six sensors setup (head, hands, hip, and feet), DragPoser surpasses AvatarPoser+ and FinalIK in pose quality, aligning with SparsePoser’s performance in Joint Error. Most importantly, DragPoser excels in End-Effector Error, outperforming both deep-learning methods and slightly improving results compared to FinalIK. This superiority is primarily attributed to the Pose Optimization process in the latent space, which utilizes the Pose Autoencoder’s Decoder for achieving deep-learning-like pose quality, while concurrently allowing for the fulfillment of hard constraints akin to traditional IK. Despite the enhancement in Positional Error, DragPoser encounters a slight increase in Rotational Error. Finally, we observe that the Velocity Error performance mirrors that of SparsePoser, with FinalIK having the best performance but with a higher standard deviation.
@@
We also experimented with further reducing the input signals to four sensors placed in the head, hands and hip. This is a highly challenging scenario for leg pose reconstruction. In this context, DragPoser exhibits a slight advantage over AvatarPoser in positional and rotational errors, however, with a slightly higher Velocity Error. It is important to note that, despite the reduced sensor count, DragPoser continues to maintain high end-effector accuracy (in this case, we take into account only the four sensors). However, when only three sensors are used, removing the hip sensor, DragPoser struggles to achieve the pose quality seen in AvatarPoser. On the other hand, when the three sensors are placed on the hip and hands, DragPoser attains comparable results to AvatarPoser when robust constraints are incorporated. We attribute this limitation in our method to the absence of a dedicated component for global position prediction. As a result, when no hip sensor is available, the method optimizes the outcome solely based on the provided sensors and the Temporal Predictor, which is proved to be insufficient for achieving accurate global positioning. To address this, we dynamically introduce losses, such as enforcing proximity of the feet to the floor and minimizing the distance between hip and head ground projections. While this adaptability highlights DragPoser’s capability to swiftly integrate new constraints, it also emphasizes the ongoing challenge of achieving precise global position accuracy without a hip sensor. In summary, our findings demonstrate that DragPoser achieves state-of-the-art performance across diverse end-effector scenarios and excels in pose reconstruction, particularly when high end-effector accuracy is crucial. In addition, DragPoser proves to be the most flexible method concerning input configuration and robustness to occlusions and sensor malfunctions, as indicated in Table 2.
@@
Figure 5. Visual comparison of the pose reconstruction results from AvatarPoser+ (AP+), FinalIK (FIK), SparsePoser (SP), and DragPoser (Ours), for seven example poses shown in the top row as ground truth (GT). All methods are reconstructed using six 6-DOF trackers as input data. We highlighted with a zoom-in the main errors found for each method when compared against the ground truth. Deep-learning-based methods (AvatarPoser+ and SparsePoser) often cannot accurately place end-effectors, while traditional IK (FinalIK) produces unnatural poses. Our method can produce high-quality poses while achieving high end-effector accuracy.
@@
We present a visual comparison in Figure 5, showcasing selected frames from different poses in the evaluation datasets. In the text, we refer to specific poses in Figure 5 as row/column. All methods utilize six 6-DoF sensors, with end-effectors represented by green spheres. Note that some methods, like FinalIK and DragPoser, use the ankle joints for foot end-effectors, while others, including AvatarPoser+ and SparsePoser, use the toe joints.
@@
Generally, FinalIK demonstrates accurate end-effector placement but often at the cost of losing plausible poses. FIK/2 and FIK/7 exemplify this, with incorrectly oriented elbows and significant deviations from ground truth in the rest of the examples. Moreover, due to its reliance on ad-hoc solutions for natural pose generation, FinalIK struggles in less common scenarios, failing to reconstruct poses or reach end-effectors, as seen in the FIK/5 and FIK/6.
@@
Both AvatarPoser+ and SparsePoser encounter similar challenges. While they are able to decode high-quality poses, there are instances, such as in the AP+/4 and AP+/6, where AvatarPoser+ fails to accurately reconstruct poses. In terms of end-effectors, most poses from these methods do not precisely align with the target end-effectors. We believe this happens with traditional deep-learning-based approaches due to the use of large networks; since high-frequency details are lost during the forward process.
@@
DragPoser consistently generates high-quality poses across all cases, successfully meeting the end-effector hard constraints. Notably, it excels in reconstructing challenging poses like push-ups Ours/6 and kicks Ours/4, effectively balancing pose accuracy with the fulfillment of hard constraints.
5.2. Dynamic Constraints Evaluation
Sensors DoF Joint Error End-Effector
Pos (cm) Rot (deg) Vel (cm/s) Pos (cm)
3 Head+Hands 6 15.5(15.0) 22.0(22.3) 34.0(50.2) 16.1(11.4)
3 Hip+Hands 6 9.95(13.7) 14.1(15.6) 23.1(36.2) 1.77(1.31)
4 Hip+Head+Hands 6 7.86(13.1) 10.6(12.4) 20.0(41.5) 1.12(0.72)
5 Hip+Head+Hands+RFoot 6 4.95(9.44) 9.25(10.8) 15.3(32.8) 1.17(0.77)
6 Hip+Head+Hands+Feet; Faulty 1% 6 6.55(11.1) 12.0(14.8) 21.1(59.7) 10.2(16.0)*
6 Hip+Head+Hands+Feet; Faulty 0.5% 6 4.50(8.50) 9.82(12.1) 17.0(50.4) 5.85(11.9)*
6 Hip+Head+Hands+Feet 3 Pos 5.81(6.71) 25.9(26.2) 13.2(19.6) 1.56(1.51)
6 Hip+Head+Hands+Feet 6 2.18(2.70) 7.66(8.71) 10.7(15.1) 1.02(0.73)
@@ 5.2. Dynamic Constraints Evaluation
Table 2. Accuracy of our method for very diverse sparse input scenarios. Each scenario is determined by the number of sensors (3 to 6), their placement, and potential faulty behavior (first column), as well as the DoFs of each sensor (3 for position-only, 6 for position+rotation; second column). We report the joint errors and the end-effector errors, as in Table 1.

### v2

@@ 5 Experiments and Evaluation
For evaluation, we trained DragPoser using the DanceDB dataset [DanceDB:2019] and evaluated with the HUMAND4D [HUMAN4D:2020] and SOMA [SOMA:2021] datasets, both part of the AMASS collection [AMASS:2019]. Our choice of AMASS was guided by its public availability and its use in previous works. Additionally, for some images and supplemental material, we employed the SparsePoser database [ponton2023sparseposer]. This database is particularly advantageous for VR applications, offering diverse common motions and reducing incorrect poses through the use of IMU-based motion capture compared to the optical systems used in AMASS datasets. The Pose Optimizer process is initialized with the first pose of each motion clip.
@@
We implemented DragPoser using PyTorch and optimized the networks with the AdamW optimizer [Loshchilov:2019]. The system was trained and evaluated on a PC configured with an Intel Core i7-12700k CPU, 32GB of RAM, and an NVIDIA GeForce RTX 3090 GPU. The Pose Autoencoder was trained with a batch size of 64 and a learning rate of $10^{-4}$ over approximately one day; we set the latent space dimension to $L=24$ and the number of joints to $J=22$, resulting in about 168 thousand parameters. For the Temporal Predictor, we trained it with a batch size of 512 and a learning rate of $10^{-3}$ for approximately thirty minutes. The Transformer hyperparameters were set to four heads, three layers each for the encoder and decoder, a feedforward dimension of 2048, and a feature dimension of 48, totaling approximately 1.2 million parameters. During training, we used $W_{p}=16$ and $W_{f}=16$ and ran the Pose Autoencoder’s Encoder to generate the training database for the Temporal Predictor. Additionally, we empirically found that introducing random noise to the four limbs aids in regularizing the Temporal Predictor. Thus, in every training iteration, we introduce Gaussian noise to the latent vector, with mean and standard deviation values for each dimension derived from the dataset, applied to each limb with a 10% probability. We initially experimented with adding noise to arbitrary joints in the full skeleton, but we found that targeting individual limbs was more effective. First we train the Pose Autoencoder with a weighted combination of all the losses presented $\mathcal{L}=\lambda_{Q}\mathcal{L}_{Q}+\lambda_{FK}\mathcal{L}_{FK}+\lambda_{KLD}\mathcal{L}_{KLD}+\lambda_{C}\mathcal{L}_{C}$ where $\lambda_{Q}=1$, $\lambda_{FK}=100$, $\lambda_{KLD}=0.001$ and $\lambda_{C}=1$. Subsequently, we train the Temporal Predictor with a standard MSE loss.
5.1 Comparison
We compare DragPoser with three state-of-the-art deep-learning-based methods, AvatarPoser [Jiang:2022b], SparsePoser [ponton2023sparseposer] and HuMoR [rempe:2021], as well as FinalIK, a state-of-the-art IK method for animating full-body characters. These methods were selected due to their public availability and their representation of different sensor configurations and approaches.
AvatarPoser utilizes a Transformer encoder to generate full-body poses from sparse input (three or four sensors) and refines the arms with an optimization-based IK. Notably, it excels in producing high-quality poses in scenarios with high ambiguity and limited sensors. We extended the method, referred to here as AvatarPoser+, by adding two additional sensors on the feet, allowing a fair comparison in six sensor setups. Note that AvatarPoser was retrained for each different experiment due to its fixed input.
In addition, we compare with SparsePoser, the state-of-the-art deep-learning-based method in terms of pose quality and end-effector accuracy for six 6-DoF sensors. Their approach is fixed to six input sensors and uses an autoencoder and learned IK networks for full-body pose reconstruction.
We also compare it with HuMoR, which, similar to our work, uses a VAE autoencoder to construct a latent space later used during an optimization process for multiple tasks. Conversely, HuMoR employs a VAE autoencoder on pose transitions rather than complete poses. Additionally, the provided implementation requires processing entire motion sequences, increasing computational demands and leaving the method unsuitable for real-time or interactive applications. Moreover, motion sequence length in HuMoR is severely constrained by available memory, requiring batch processing and additional techniques to maintain continuity. For this reason, we executed HuMoR in batches of 60 frames as recommended in the original paper since we experimented with longer batches, but the resulting animations were often highly inaccurate.
AvatarPoser, SparsePoser and HuMoR were re-trained using the DanceDB dataset following the training procedure of their official implementations. Finally, considering the findings by Ponton et al. [ponton2023sparseposer] that FinalIK exhibits superior end-effector accuracy compared to deep learning-based methods, it is employed as a baseline for evaluating hard constraints.
Table 1: Accuracy of the reconstructed poses compared against state-of-the-art methods, for diverse sensor configurations. We report the mean and standard deviation (within parenthesis) of joint errors (position, rotation, and velocity) as well as the end-effector positional error, all of them with respect to ground truth data. Please refer to the text for details about training and test datasets. When using 6 sensors (placed at the hip, head, hands, and feet), we compare against AvatarPoser+ (i.e., [Jiang:2022] extended to use also the two sensors on the feet), FinalIK, HuMoR [rempe:2021] and SparsePoser [ponton2023sparseposer]. For the 4-sensor configuration (hip, head, and hands), we compare against AvatarPoser and HuMoR; FinalIK is not considered since it requires all end-effectors, and SparsePoser’s architecture is designed for a fixed 6-sensor configuration. Concerning the 3-sensor setup, we tested our approach with the head+hands and hip+hands configurations, and compared it against the default head+hands AvatarPoser configuration and HuMoR.
HuMoR reconstructs motion in sequences, allowing it to leverage future information to adjust past poses and, thus, refine entire motion sequences during offline processing—in contrast to the other real-time, frame-to-frame prediction methods used in the comparison.
Method Sensors End-Effect. Err. Joint Error
@@ 5.1 Comparison
AvatarPoser+ 6 11.2(7.93) 11.3(10.5) 13.4(28.5) 7.66(7.60)
FinalIK 6 1.16(1.22) 11.5(19.2) 6.18(21.2) 3.31(4.64)
HuMoR 6 9.25(4.91) 9.35(10.4) 19.8(35.9) 6.96(5.27)
SparsePoser 6 3.43(2.53) 5.78(5.85) 10.8(15.0) 2.81(2.66)
Ours 6 1.02(0.73) 7.66(8.71) 10.7(15.1) 2.18(2.70)
AvatarPoser 4 9.62(6.51) 12.4(11.7) 16.3(31.7) 9.24(10.1)
HuMoR 4 8.61(4.14) 10.8(11.3) 25.6(56.4) 10.9(11.6)
Ours 4 1.12(0.72) 10.6(12.4) 20.0(41.5) 7.86(13.1)
AvatarPoser 3 12.3(7.51) 14.3(12.8) 20.1(39.5) 11.0(10.8)
Head+Hands
HuMoR 3 15.0(8.30) 11.3(12.5) 28.0(65.1) 14.2(13.0)
Head+Hands
HuMoR 3 9.92(4.71) 11.6(12.6) 27.6(63.0) 13.4(12.8)
Hip+Hands
Ours 3 16.1(11.4) 22.0(22.3) 34.0(50.2) 15.5(15.0)
Head+Hands
Ours 3 1.77(1.31) 14.1(15.6) 23.1(36.2) 9.95(13.7)
Hip+Hands
@@
In the six sensors setup (head, hands, hip, and feet), DragPoser surpasses AvatarPoser+, FinalIK and HuMoR in pose quality, aligning with SparsePoser’s performance in Joint Error. Most importantly, DragPoser excels in End-Effector Error, outperforming all deep-learning methods and slightly improving results compared to FinalIK. This superiority is primarily attributed to the Pose Optimization process in the latent space, which utilizes the Pose Autoencoder’s Decoder for achieving deep-learning-like pose quality, while concurrently allowing for the fulfillment of hard constraints as traditional IK. Despite the enhancement in Positional Error, DragPoser encounters a slight increase in Rotational Error. Finally, we observe that the Velocity Error performance mirrors that of SparsePoser, with FinalIK having the best performance but with a higher standard deviation.
We also experimented with further reducing the input signals to four sensors placed in the head, hands and hip. This is a highly challenging scenario for leg pose reconstruction. In this context, DragPoser exhibits a slight advantage over AvatarPoser and HuMoR in positional and rotational errors, however, with a slightly higher Velocity Error compared to AvatarPoser. It is important to note that, despite the reduced sensor count, DragPoser continues to maintain high end-effector accuracy (in this case, we take into account only the four sensors).
@@
However, when only three sensors are used, and the hip sensor is removed, DragPoser struggles to achieve the pose quality seen in AvatarPoser and HuMoR. On the other hand, when the three sensors are placed on the hip and hands, DragPoser attains comparable results to AvatarPoser and HuMoR when robust constraints are incorporated. We attribute this limitation in our method to the absence of a dedicated component for global position prediction in the case of AvatarPoser, and not optimizing at a sequence level such as HuMoR. However, we decided to prioritize frame-to-frame prediction to use our method in an online prediction setting. As a result, when no hip sensor is available, the method optimizes the outcome solely based on the provided sensors and the Temporal Predictor, which proved insufficient for achieving accurate global positioning. To address this, we dynamically introduce losses, such as enforcing proximity of the feet to the floor and minimizing the distance between hip and head ground projections. While this adaptability highlights DragPoser’s capability to swiftly integrate new constraints, it also emphasizes the ongoing challenge of achieving precise global position accuracy without a hip sensor.
In summary, our findings demonstrate that DragPoser achieves state-of-the-art performance across diverse end-effector scenarios and excels in pose reconstruction, particularly when high end-effector accuracy is crucial. In addition, DragPoser proves to be a flexible method concerning input configuration and robustness to occlusions and sensor malfunctions, as indicated in Table 2.
@@
Figure 4: Visual comparison of the pose reconstruction results from HuMoR, AvatarPoser+ (AP+), FinalIK (FIK), SparsePoser (SP), and DragPoser (Ours). The ground truth poses are highlighted in green and overlaid on the results. All methods are reconstructed using six 6-DOF trackers as input data. We highlighted with a zoom-in some examples of errors found when compared against the ground truth. Deep-learning-based methods (HuMoR, AvatarPoser+ and SparsePoser) often cannot accurately place end-effectors, while traditional IK (FinalIK) produces unnatural poses. Our method can produce high-quality poses while achieving high end-effector accuracy.
@@
We present a visual comparison in Figure 4, showcasing selected frames from different poses in the evaluation datasets. In the text, we refer to specific poses in Figure 4 as row/column. All methods utilize six 6-DoF sensors, with end-effectors represented by green spheres. Note that some methods—HuMoR, FinalIK and DragPoser—use the ankle joints for foot end-effectors, while the others—AvatarPoser+ and SparsePoser—use the toe joints.
@@
Generally, FinalIK demonstrates accurate end-effector placement but often at the cost of losing plausible poses. FIK/1 and FIK/3 exemplify this, with incorrectly oriented elbows and significant deviations from ground truth in the rest of the examples. Moreover, due to its reliance on ad-hoc solutions for natural pose generation, FinalIK struggles in less common scenarios, failing to reconstruct poses or reach end-effectors, as seen in the FIK/4 and FIK/5.
@@
HuMoR, AvatarPoser+ and SparsePoser encounter similar challenges. While they are able to decode high-quality poses, there are instances, such as in the AP+/3, AP+/5 or HuMoR/5, where they fail to accurately reconstruct poses. In terms of end-effectors, most poses from these methods do not precisely align with the target end-effectors. We believe this happens with traditional deep-learning-based approaches due to the use of large networks; since high-frequency details are lost during the forward process.
We observed that latent space optimization methods like HuMoR often encounter difficulties when generating poses for long sequences (e.g., $>60$ poses). This is attributed to the structure and navigation of the latent space, which can lead to the traversal of invalid regions and, subsequently, incorrect results (see our supplemental material for more details). DragPoser mitigates these issues through frame-by-frame optimization, ensuring the continuity of the latent space structure, and utilizing a Temporal Predictor to guide the optimization process.
@@
DragPoser consistently generates high-quality poses across all cases, successfully meeting the end-effector hard constraints. Notably, it excels in reconstructing challenging poses like push-ups Ours/5 and kicks Ours/3, effectively balancing pose accuracy with the fulfillment of hard constraints.
5.2 Dynamic Constraints Evaluation
@@
Table 2: Accuracy of our method for very diverse sparse input scenarios. Each scenario is determined by the number of sensors (3 to 6), their placement, and potential faulty behavior (first column). All sensors have 6 DoFs (position+rotation) unless otherwise noted. We report the joint errors and the end-effector errors, as in Table 1.
*Reported end-effector errors encompass all six sensors; this error is significantly larger compared to more stable setups where only the tracked sensors are evaluated.
Sensors End-Effector Joint Error
Pos (cm) Pos (cm) Rot (deg) Vel (cm/s)
3 Head+Hands 16.1(11.4) 22.0(22.3) 34.0(50.2) 15.5(15.0)
3 Hip+Hands 1.77(1.31) 14.1(15.6) 23.1(36.2) 9.95(13.7)
4 Hip+Head+Hands 1.12(0.72) 10.6(12.4) 20.0(41.5) 7.86(13.1)
5 Hip+Head+Hands+RFoot 1.17(0.77) 9.25(10.8) 15.3(32.8) 4.95(9.44)
6 Hip+Head+Hands+Feet 10.2(16.0)* 12.0(14.8) 21.1(59.7) 6.55(11.1)
Faulty 1%
6 Hip+Head+Hands+Feet 5.85(11.9)* 9.82(12.1) 17.0(50.4) 4.50(8.50)
Faulty 0.5%
6 Hip+Head+Hands+Feet 1.56(1.51) 25.9(26.2) 13.2(19.6) 5.81(6.71)
3 DoF Position
6 Hip+Head+Hands+Feet 1.02(0.73) 7.66(8.71) 10.7(15.1) 2.18(2.70)